# ref.bib UUITP-22/24

### v1

@@ 2.2 Equivariant disk potential of generic toric threefolds
where $[\mathcal{H}^{D}]_{0}$ is defined as the order zero term in $\xi$ in the series expansion of $\mathcal{H}^{D}(t,c,\xi\epsilon)$, analogously to (2.19), and the ellipses denote terms that are polynomial in $c$, which therefore do not contribute to the disk potential.
Plugging this formula in (2.26) demonstrates that the regular part of the equivariant superpotential actually computes the disk potential $W_{\mathrm{disk}}$. In fact, the latter is defined by its relation to the roots of the mirror curve curve as follows
$\partial_{c}W^{(j)}_{\mathrm{disk}}=\log(-y_{j})+\dots\,,$ (2.29)
for a brane whose vacuum configuration is labeled by the $j$-th sheet of the mirror curve. Equation (2.28) is actually a generalization of the claim in (2.20), because instead of a single disk potential it features a linear combination of them, weighted by the asymptotic slopes $\alpha_{j}$. Note however, that the disk instantons $W^{(j)}_{\mathrm{disk}}$ on different branches are actually related, since roots $y_{j}(x)$ are related to each other by analytic continuation around branch points of the mirror curve. [^7] For example, if the mirror curve is quadratic in $y$ then $H(x,y)=y^{2}+a_{1}(x)y+a_{2}(x)$ where $a_{2}(x)=y_{1}y_{2}$ is the product of the two roots. Equation (2.29) then implies that $W_{\mathrm{disk}}^{(2)}=-W_{\mathrm{disk}}^{(1)}+\int^{c}\,\log a_{2}(x)\,\mathop{}\!\mathrm{d}\log x$. This explains why in all examples considered in [Cassia:2023uvd], the simpler statement (2.20) holds, with suitable proportionality coefficients.
An remark on the validity of formula (2.28) is in order. Its derivation makes use of a series expansion of $F_{D}^{(n)}$ in (2.24) whose domain of convergence is the product of unit disks
$|1+y_{j}^{-1}|<1\qquad j=1,\dots,k\,.$ (2.30)
In particular, this condition must hold for the asymptotic values of the roots $y_{j}(x)$ of (2.23) within some region of the large volume phase, which is $x\to 0$ in this case. From 2.27, it follows that this condition can be satisfied if e.g. all $\alpha_{j}>0$ or if one of them, say $\alpha_{1}$ is zero with $\lim_{x\to 0}y_{1}=-1$. The latter case, which features a ‘horizontal’ asymptotic branch of the mirror curve, is conventionally adopted in computations of open string instantons as in [Aganagic:2000gs, Aganagic:2001nx].
In most models the condition can be satisfied by choosing coordinates for $H(x,y)$ in such a way that all $y_{j}$ are large or constant. For example in the phase $x\to 0$ this means that $\alpha_{j}\geq 0$ for all $j$. If this is the case, it is generically possible to find a sector around $x=0$ where (2.30) is satisfied. However, even this condition on $\alpha_{j}$ is not always sufficient, was will be seen in examples.
Remark 2.4.
In general, condition (2.30) is not always true. However, this does not invalidate the relation between monodromy of $\mathcal{H}^{D}$ and $W_{\mathrm{disk}}$. Instead, when this condition is violated it simply means that one needs to find another way to expand the Lauricella hypergeometric. It is generally possible to compute series expansions of $F_{D}^{(n)}$ in other regions of moduli space by means of analytic continuation formulas, see e.g. [Bezrodnykh_2018] and references therein for examples. In fact, in [Cassia:2023uvd] we addressed one such example in the case of local $\mathbb{P}^{2}$, where the equivariant superpotential was computed in a phase in which (2.30) is violated. [^8] In that instance the Lauricella function specializes to simpler hypergeometric series, whose analytic continuations are known.
@@ 2.3.1 $\mathbb{C}^{3}$
First, we check that the general expression of $\mathcal{H}^{D}$ in terms of Lauricella’s function reproduces the quantum Lebesgue measure, which is known to be (see [Cassia:2023uvd, (3.11)] for a derivation)
@@
In order to match this expression with the general formula in (2.24), we write $\mathcal{H}^{D}$ as a $\epsilon$-GLSM hemisphere partition function and then rewrite each $\operatorname{\Gamma}$ in the integrand using
@@
from which we read
@@ 2.3.2 Local $\mathbb{P}^{2}$
For this choice of cut and assuming $t>0$, there exist three distinct phases for the associated toric Lagrangian: $c>0$, $0>c>-t$ and $-t>c$, respectively. The quantum Lebesgue measure was computed in [Cassia:2023uvd, eq. (5.23)] in the phase $0>-t>c$
@@
The same expression holds in other phases, by analytic continuation. This should be compared to the general formula (2.24). We repeat the derivation for clarity and to match variables, starting from the $\epsilon$-GLSM integral encoded by (2.40)
@@
In particular, this expression makes it clear that if one of the arguments lies within the unit disk, the other must lie outside. Therefore this is an example in which condition (2.30) is not satisfied, and formula (2.28) is not expected to hold. 101010 Indeed, it is easy to see that applying this formula would lead to a wrong result. Since $y_{1}y_{2}=zx^{3}$ the first term in (2.28) would not give any interesting contribution, while the second term would vanish since $\alpha_{1}=\alpha_{2}$ and because of this simple relation between the roots. We emphasize that the expression (2.24) of $\mathcal{H}^{D}$ in terms of Lauricella’s function still holds. Its monodromy however needs to be computed by other means, such as those adopted in [Cassia:2023uvd], or by relying on known analytic continuation formulae (see e.g. [Bezrodnykh_2018] and references therein).
@@ 3.1 Braverman’s construction
Using the restriction $z_{+}z_{-}=w$ into the moment map condition gives $\mu_{q}(x)-c=|z_{+}|^{2}-|w|^{2}/|z_{+}|^{2}$. Solving for $|z_{+}|^{2}$ and choosing the (globally) positive root, defines a circle in the $\mathbb{C}_{+}$ plane, for each $x\in X_{3}$. The circle has nonzero radius thanks to the condition $w\neq 0$ which implies $z_{\pm}\neq 0$, therefore the moment map equation combined with the restriction to the fibre gives a manifold with topology $X_{3}\times S^{1}$. Taking the $U(1)$ quotient reduces the circle to a point, showing that to each $w$ there corresponds a whole copy of $X_{3}$. Therefore we conclude that the generic fibre of $\pi$ is complex isomorphic to $X_{3}$
@@
At $w=0$ the fibre is different. In this case there are two components to consider, corresponding to $z_{\pm}=0$ and $z_{\mp}\in\mathbb{C}_{\mp}$ which intersect at the origin of $\mathbb{C}_{+}\times\mathbb{C}_{-}$. Each leads to a reduction of (3.3) to a 3-manifold that, by definition, corresponds to one of the ‘half spaces’ of the symplectic cut:
@@
Finally, we obtain that the fibre at $w=0$ is
@@
To summarize, Braverman’s construction defines a Calabi–Yau 4-fold $X_{4}$ which admits the structure of a fibration over a complex plane $\mathbb{C}$, whose generic fiber is a smooth Kähler manifold complex isomorphic to $X_{3}$ and with a singular fiber at the origin which is complex isomorphic to $X_{3}^{<}\cup_{X_{2}}X_{3}^{>}$. [^12] The isomorphism as complex manifolds is equivariant with respect to the $S^{1}$-action. The generic fibre is not isomorphic to $X_{3}$ as a Kähler manifold, see [Braverman+1999+85+98, Remark 3.4]. This fibration therefore realizes the symplectic cut as a degeneration of the Calabi–Yau threefold $X_{3}$. The half-spaces of the symplectic cut are complex isomorphic to toric divisors of $X_{4}$, as made explicit in (3.8) and (3.9). Correspondingly, the common submanifold $X_{2}$ arises as the intersection of these divisors as shown by (3.10). The global picture is summarized by Figure 2.
@@ 3.2 Quantization of Braverman’s fourfold
Recall that $X_{4}$ is a fibration of $X_{3}$ over $\mathbb{C}$ (as in (3.5)), and that $X_{3}$ itself is sliced by $X_{2}$ (as in (2.11)). We therefore expect to be able to decompose the quantum volume $\mathcal{F}^{D}_{X_{4}}$ in terms of the function $\mathcal{H}^{D}$. In order to see this explicitly, we introduce the Fourier transform [^13] This is defined for $\varphi\in\mathrm{i}\mathbb{R}$ and extended by analytic continuation. of the quantum volume of $X_{2}$
@@
$=$ $\operatorname{\Gamma}(\epsilon_{+}+\epsilon_{-})\int\frac{\mathop{}\!\mathrm{d}\varphi}{2\pi\mathrm{i}}\,\mathrm{e}^{\varphi c}\,B(\epsilon_{+}-\varphi,\epsilon_{-}+\varphi)$
@@
Note that this is the quantum volume of a onefold obtained by taking the symplectic quotient +×$\mathbb{C}\cong\mathbb{C}_{+}\times\mathbb{C}_{-}\sslash U(1)$ with charges $(-1,+1)$, and it is a multi-valued function with ramification at $\mathrm{e}^{-c}\in\{0,-1,\infty\}$ therefore the computation of the monodromy must be handled with some care. This also gives a geometric interpretation to formula (3.16), as the computation of a 4d quantum volume in terms of a fibration over the complex line by a product of a twofold with quantum volume $\mathcal{H}^{D}$ and a onefold with quantum volume $\rho$.
Next, recall that from the viewpoint of $X_{4}$ the half-spaces $X_{3}^{\lessgtr}$ arise as divisors, as described in (3.8) and (3.9). In the framework of $\epsilon$-GLSM, equivariant periods of divisors, by which we mean solutions to the PF equations of $X_{4}$ with classical behavior equal to the (equivariant) volume of the corresponding divisor, can be obtained from $\mathcal{F}^{D}_{X_{4}}$ by acting with difference operators as
@@ 3.3 Quantum cuts of CY3 from equivariant periods of CY4
Observe that for small ϵ±subscriptitalic-$\epsilon_{\pm}$, Euler’s beta function has leading order behaviour141414 This is a well-known property, that can be derived using the integral representation $B(a,b)=\int_{0}^{\infty}\frac{t^{a-1}}{(1+t)^{a+b}}\,\mathop{}\!\mathrm{d}t$ as follows $B(\mathrm{i}x,-\mathrm{i}x)=\int_{0}^{\infty}t^{\mathrm{i}x-1}\,\mathop{}\!\mathrm{d}t=\int_{-\infty}^{\infty}\mathrm{e}^{\mathrm{i}xy}\,\mathop{}\!\mathrm{d}y=2\pi\,\delta(x)\,.$ (3.26)
$B(\epsilon_{+}-\phi,\epsilon_{-}+\phi)=B(-\phi,\phi)\left(1+O(\epsilon_{\pm})\right)=2\pi\mathrm{i}\,\delta(\phi)\left(1+O(\epsilon_{\pm})\right)$ (3.27)
@@
This leads to the conclusion that, in the limit $\epsilon_{\pm}\to 0$, the period associated to $X_{2}$ computed as intersection of divisors $D_{\pm}$ in $X_{4}$, coincides with the equivariant disk potential computed from the quantum cut of $X_{3}$. Moreover, the derivation of the limit in (3.34) also gives an alternative integral expression for the equivariant superpotential as
@@
Remark 3.2 (Quantum volume of $X_{2}$ vs equivariant period of divisor intersection).
@@ 3.4 Extended Picard–Fuchs equations for quantum cuts
This is a direct consequence of the observation that both $\mathcal{F}^{D}_{\lessgtr}$, as well as the equivariant superpotential $W$, arise from the limit $\epsilon_{\pm}\to 0$ of equivariant periods of divisors (and their intersection) in $X_{4}$.
@@
Taking the limit $\epsilon_{\pm}\to 0$ of these equations leads to Picard–Fuchs equations involving open string moduli for the equivariant superpotential and for the half-volumes defined by the quantum cut of $X_{3}$, namely
@@
We thus find that half-volumes and the equivariant superpotential obey an extension of the Picard–Fuchs equations, where $\operatorname{\mathcal{D}}_{i}$ are modified by the addition of $q_{i}\partial_{c}$. Of course, these equations also admit solutions that are independent of $c$, which correspond to $\mathcal{F}^{D}$ and equivariant periods of intersections of divisors of $X_{3}$. However, when considering more general solutions that do depend on $c$, we find among these $\mathcal{F}^{D}_{\lessgtr}$ and $W$.
@@
Remark 3.4 (Relation to work of Mayr and Lerche).
In [Mayr:2001xk, Lerche:2001cw] Mayr and Lerche observed that the disk potential of open topological strings in toric Calabi–Yau threefolds with a toric brane arise from periods of certain Calabi–Yau fourfolds. Our construction of the equivariant superpotential and the observation, based on Braverman’s construction, that it corresponds to the limit $\epsilon_{\pm}\to 0$ of an intersection of divisors in $X_{4}$, provides an alternative derivation of the results of Mayr–Lerche. Our construction generalizes this to the equivariant setting, and also makes is clear that the open-closed string correspondence between threefolds and fourfolds works for toric branes in arbitrary toric Calabi–Yau threefolds.

### v2

@@ 2.2 Equivariant disk potential of general toric threefolds
where the ellipses denote terms that are polynomial in $c$, which therefore do not contribute to the disk potential.
Because of identity (2.22), the expression we just derived also gives the regular instantonic part of $\partial_{c}W(t,c,\epsilon)$, so it remains now to show that this matches with the disk potential as in (2.27). This can be achieved by observing that $W_{\mathrm{disk}}$ is in fact defined by its relation to the roots of the mirror curve as follows
$\partial_{c}W^{(j)}_{\mathrm{disk}}=\log(-y_{j}(-x))+\dots\,,$ (2.33)
for a brane whose vacuum configuration is labeled by the $j$-th sheet of the mirror curve. This is precisely the type of logarithmic contributions we have found in the r.h.s. of (2.32). Interestingly, we can conclude that equation (2.32) actually leads to a generalization of both claims in (2.26) and (2.27), because instead of a single disk potential it features a linear combination of them, weighted by the asymptotic slopes $\alpha_{j}$. Note however, that the disk instantons $W^{(j)}_{\mathrm{disk}}$ on different branches are actually related, since roots $y_{j}(x)$ are related to each other by analytic continuation around branch points of the mirror curve. [^9] For example, if the mirror curve is quadratic in $y$ then $H(x,y)=y^{2}+a_{1}(x)y+a_{2}(x)$ where $a_{2}(x)=y_{1}y_{2}$ is the product of the two roots. Equation (2.33) then implies that $W_{\mathrm{disk}}^{(2)}=-W_{\mathrm{disk}}^{(1)}+\int^{c}\,\log a_{2}(x)\,\mathop{}\!\mathrm{d}\log x$. This explains why in all examples considered in [Cassia:2023uvd], the simpler statement (2.26) holds, with suitable proportionality coefficients.
@@ 2.3.1 $\mathbb{C}^{3}$
First, we check that the general expression of $\mathcal{H}^{D}$ in terms of Lauricella’s function reproduces the quantum Lebesgue measure, which was shown to be (see [Cassia:2023uvd, (3.11)] for a derivation)
@@
To exemplify the general approach outlined in the last section, we are now going to show how to match (2.35) with the general formula in (2.30). In order to do so, we start from the definition of $\mathcal{H}^{D}$ as a $\epsilon$-GLSM hemisphere partition function as in (2.17) and then rewrite each $\operatorname{\Gamma}$ in the integrand using
@@
where in the second line we have obtained a Dirac $\delta$-function after integrating over $\varphi$, and next we used a reparametrization of the integration variables [^10] The choice of which variable $u_{i}$ to rescale by and integrate out first is not canonical and different choices lead to different but equivalent expressions for $\mathcal{H}^{D}$. Similar manipulations can be performed in the general context of an arbitrary single or double cut of a CY threefold as discussed in Appendix B $u_{1,2}=u_{3}s_{1,2}$ and later the redefinition $s_{1}=y$, $s_{2}=x$. From this expression of $\mathcal{H}^{D}$ we can read
@@ 2.3.2 Local $\mathbb{P}^{2}$
For this choice of cut and assuming $t>0$, there exist three distinct phases for the associated toric Lagrangian: $c>0$, $0>c>-t$ and $-t>c$, respectively. The quantum volume of local $\mathbb{P}^{2}$ before the cut is given by the integral
$\mathcal{F}^{D}(t,\epsilon)=\int_{\mathrm{i}\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\phi}{2\pi\mathrm{i}}\mathrm{e}^{t\phi}\operatorname{\Gamma}(\epsilon_{1}+\phi)\operatorname{\Gamma}(\epsilon_{2}+\phi)\operatorname{\Gamma}(\epsilon_{3}+\phi)\operatorname{\Gamma}(\epsilon_{4}-3\phi)$ (2.44)
which was evaluated explicitly in [Cassia:2023uvd, eq. (C.3)]. Similarly, the quantum Lebesgue measure was computed in [Cassia:2023uvd, eq. (5.23)] in the phase $0>-t>c$
@@
The same expression holds in other phases, by analytic continuation.
Remark 2.2.
This should be compared to the general formula (2.30). We repeat the derivation for clarity and to match variables, starting from the $\epsilon$-GLSM integral encoded by (2.43)
@@
In particular, this expression makes it clear that if one of the arguments lies within the unit disk, the other must lie outside. Therefore this is an example in which condition (C.19) is not satisfied, and formula (2.32) is not expected to hold. 121212 Indeed, it is easy to see that applying this formula would lead to a wrong result. Since $y_{1}y_{2}=zx^{3}$ the first term in (2.32) would not give any interesting contribution, while the second term would vanish since $\alpha_{1}=\alpha_{2}$ and because of this simple relation between the roots. We emphasize that the expression (2.30) of $\mathcal{H}^{D}$ in terms of Lauricella’s function still holds. Its monodromy however needs to be computed by other means, such as those adopted in [Cassia:2023uvd], or by relying on known analytic continuation formulae (see e.g. [Bezrodnykh_2018] and references therein).
@@ 3.1 Braverman’s construction
Inserting the restriction $z_{+}z_{-}=w$ into the moment map condition gives $\mu_{q}(x)-c=|z_{+}|^{2}-|w|^{2}/|z_{+}|^{2}$. Solving for $|z_{+}|^{2}$ and choosing the (globally) positive root, defines a circle in the $\mathbb{C}_{+}$ plane, for each $x\in X_{3}$. The circle has nonzero radius thanks to the condition $w\neq 0$ which implies $z_{\pm}\neq 0$, therefore the moment map equation combined with the restriction to the fiber gives a manifold with topology $X_{3}\times S^{1}$. Taking the $U(1)$ quotient reduces the circle to a point, showing that to each $w$ there corresponds a whole copy of $X_{3}$. Therefore we conclude that the generic fiber of $\pi$ is complex isomorphic to $X_{3}$
@@
At $w=0$ the fiber is different. In this case there are two components to consider, corresponding to $z_{\pm}=0$ and $z_{\mp}\in\mathbb{C}_{\mp}$ which intersect at the origin of $\mathbb{C}_{+}\times\mathbb{C}_{-}$. Each leads to a reduction of (3.3) to a 3-manifold that, by definition, corresponds to one of the ‘half spaces’ of the symplectic cut:
@@
Finally, we obtain that the fiber at $w=0$ is
@@
To summarize, Braverman’s construction defines a Calabi–Yau fourfold $X_{4}$ which admits the structure of a fibration over a complex plane $\mathbb{C}$, whose generic fiber is a smooth Kähler manifold complex isomorphic to $X_{3}$ and with a singular fiber at the origin which is complex isomorphic to $X_{3}^{<}\cup_{X_{2}}X_{3}^{>}$. [^14] The isomorphism as complex manifolds is equivariant with respect to the $S^{1}$-action. The generic fiber is not isomorphic to $X_{3}$ as a Kähler manifold, see [Braverman+1999+85+98, Remark 3.4]. This fibration therefore realizes the symplectic cut as a degeneration of the Calabi–Yau threefold $X_{3}$. The half-spaces of the symplectic cut are complex isomorphic to toric divisors of $X_{4}$, as made explicit in (3.8) and (3.9). Correspondingly, the common submanifold $X_{2}$ arises as the intersection of these divisors as shown by (3.10). The global picture is summarized by Figure 2.
@@ 3.2 Quantization of Braverman’s fourfold
Recall that $X_{4}$ is a fibration of $X_{3}$ over $\mathbb{C}$, except at the origin in $\mathbb{C}$ where the fiber $X_{3}$ degenerates to a singular space as explained in (3.5), and that $X_{3}$ itself is sliced by $X_{2}$ (as in (2.14)). Since the locus where the generic fiber becomes singular is measure zero within $X_{4}$, we expect to be able to decompose the quantum volume $\mathcal{F}^{D}_{X_{4}}$ in terms of the function $\mathcal{H}^{D}$. In order to see this explicitly, we introduce the Fourier transform [^15] This is defined for $\varphi\in\mathrm{i}\mathbb{R}$ and extended by analytic continuation. of the quantum volume of $X_{2}$
@@
$=$ $\operatorname{\Gamma}(\epsilon_{+}+\epsilon_{-})\int\frac{\mathop{}\!\mathrm{d}\varphi}{2\pi\mathrm{i}}\,\mathrm{e}^{\varphi c}\,\mathrm{B}(\epsilon_{+}-\varphi,\epsilon_{-}+\varphi)$
@@
Note that this is the quantum volume of a onefold obtained by taking the symplectic quotient +×$\mathbb{C}\cong\mathbb{C}_{+}\times\mathbb{C}_{-}\sslash U(1)$ with charges $(-1,+1)$, and it is a multivalued function with ramification at $\mathrm{e}^{-c}\in\{0,-1,\infty\}$ therefore the computation of the monodromy must be handled with some care. This also gives a geometric interpretation to formula (3.16), as the computation of a 4d quantum volume in terms of a fibration over the complex line by a product of a twofold with quantum volume $\mathcal{H}^{D}$ and a onefold with quantum volume $\rho$.
Next, recall that from the viewpoint of $X_{4}$ the half-spaces $X_{3}^{\lessgtr}$ arise as divisors, as described in (3.8) and (3.9). In the framework of $\epsilon$-GLSM, equivariant periods of divisors, by which we mean solutions to the PF equations of $X_{4}$ with classical behavior equal to the (equivariant) volume of the corresponding divisor, can be obtained from $\mathcal{F}^{D}_{X_{4}}$ by acting with difference operators as in (2.10), namely
@@ 3.3 Quantum cuts of CY3 from equivariant periods of CY4
Observe that for small ϵ±subscriptitalic-$\epsilon_{\pm}$, Euler’s beta function has leading order behavior161616 This is a well-known property, that can be derived using the integral representation $\mathrm{B}(a,b)=\int_{0}^{\infty}\frac{t^{a-1}}{(1+t)^{a+b}}\,\mathop{}\!\mathrm{d}t$ as follows $\mathrm{B}(\mathrm{i}x,-\mathrm{i}x)=\int_{0}^{\infty}t^{\mathrm{i}x-1}\,\mathop{}\!\mathrm{d}t=\int_{-\infty}^{\infty}\mathrm{e}^{\mathrm{i}xy}\,\mathop{}\!\mathrm{d}y=2\pi\,\delta(x)\,.$ (3.26)
$\mathrm{B}(\epsilon_{+}-\phi,\epsilon_{-}+\phi)=\mathrm{B}(-\phi,\phi)\left(1+O(\epsilon_{\pm})\right)=2\pi\mathrm{i}\,\delta(\phi)\left(1+O(\epsilon_{\pm})\right)$ (3.27)
@@
This leads to the conclusion that, in the limit $\epsilon_{\pm}\to 0$, the period associated to $X_{2}$ computed as intersection of divisors $D_{\pm}$ in $X_{4}$, coincides with the equivariant disk potential computed from the quantum cut of $X_{3}$. Moreover, the derivation of the limit in (3.34) also gives an alternative integral expression for the equivariant disk potential as
@@ 3.4 Extended Picard–Fuchs equations for quantum cuts
This is a direct consequence of the observation that both $\mathcal{F}^{D}_{\lessgtr}$, as well as the equivariant disk potential $W$, arise from the limit $\epsilon_{\pm}\to 0$ of equivariant periods of divisors (and their intersection) in $X_{4}$.
@@
Taking the limit $\epsilon_{\pm}\to 0$ of these equations leads to Picard–Fuchs equations involving open string moduli for the equivariant disk potential and for the half-volumes defined by the quantum cut of $X_{3}$, namely
@@
We thus find that half-volumes and the equivariant disk potential obey an extension of the Picard–Fuchs equations, where $\operatorname{\mathcal{D}}_{i}$ are modified by the addition of $q_{i}\partial_{c}$. Of course, these equations also admit solutions that are independent of $c$, which correspond to $\mathcal{F}^{D}$ and equivariant periods of intersections of divisors of $X_{3}$. However, when considering more general solutions that do depend on $c$, we find among these $\mathcal{F}^{D}_{\lessgtr}$ and $W$.
@@
3.4.1 Relation to work of Mayr and Lerche
Mayr and Lerche [Mayr:2001xk, Lerche:2001cw] observed that the disk potentials of open topological strings in toric Calabi–Yau threefolds with a toric brane can be derived from the periods of certain CY fourfolds. Liu and Yu [Liu:2021eeb, Liu:2022swc] later formalized this observation for more general CY manifolds and orbifolds. In all explicit cases we examined, the CY fourfold constructed via Braverman’s fibration matches that of Mayr and Lerche.
Our construction of the equivariant disk potential, along with our observation that it corresponds to the $\epsilon_{\pm}\to 0$ limit of the equivariant period associated with an intersection of divisors in $X_{4}$, provides an alternative formalization of the Mayr–Lerche observation for general CY3s and embedded toric branes. In particular, our construction of the fourfold $X_{4}$ is well-suited for studying equivariant Picard–Fuchs equations and quantum cohomology rings using the language of $\epsilon$-GLSMs.
Although certain aspects of Liu and Yu’s fourfold construction resemble elements of Braverman’s approach [^19] For instance, the threefold $Y$ in [Liu:2021eeb], constructed as a partial compactification of $X_{3}$, appears closely related to one of the two half-spaces $X_{3}^{\lessgtr}$. Similarly, the divisor $D$ in $Y$ seems related to the divisor $X_{2}$ in $X_{3}^{\lessgtr}$., it remains unclear whether both methods necessarily lead to the same CY fourfold $X_{4}$.